Giving that mean=34 median=32 and standard deviation as 12. Calculate the Pearson’s coefficient of skewness.
a) 0.1
b) 0.4
c) 0.5
d) 1.6
1
Expert's answer
2015-03-26T10:31:38-0400
The Pearson’s coefficient of skewness is calculated by the following formula: Sk=3*(m-md)/sd=3*(34-32)/12=3*2/12=0.5, where m is the mean, md is the median, sd is the standard deviation. Answer: c) 0.5.